40m QIL Cryo_Lab CTN SUS_Lab CAML OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log, Page 347 of 349  Not logged in ELOG logo
ID Date Author Type Category Subject
  146   Fri Nov 30 13:46:50 2007 robConfigurationElectronicsETMX oplev dead again

Quote:
In the interests of getting the Xarm alignment script working again, I reset the local damping gains for the test masses to their previous known working values (1), then I noticed that the ETMX oplev was dead. Since the scripts use the oplev motion as a readback for the optic motion, this means the script was basically blindly swinging the optics around. Some monkeying around with swapping HeNe power supplies eventually led to the conclusion that the power strip is funky, since the laser works when plugged into another power strip. Even weirder, the HeNe and the power supply indicator light have some sort of XOR relationship going on. When one works, the other doesn't. Steve will sort out this confusion later; we're good for now.


Ech. The HeNe quit again. Let's replace it and see what happens.
  145   Fri Nov 30 11:44:57 2007 robConfigurationElectronicsETMX oplev
In the interests of getting the Xarm alignment script working again, I reset the local damping gains for the test masses to their previous known working values (1), then I noticed that the ETMX oplev was dead. Since the scripts use the oplev motion as a readback for the optic motion, this means the script was basically blindly swinging the optics around. Some monkeying around with swapping HeNe power supplies eventually led to the conclusion that the power strip is funky, since the laser works when plugged into another power strip. Even weirder, the HeNe and the power supply indicator light have some sort of XOR relationship going on. When one works, the other doesn't. Steve will sort out this confusion later; we're good for now.
  144   Fri Nov 30 11:22:22 2007 ajwSummaryCDSGEO DV => LIGO DV

Quote:
Martin Hewitson of GEO600 fame has modified the cool GEO DV
to work with the LIGO NDS system with some NDS advice from Rolf (who's over in Germany this week).

I've moved it onto the 40m CDS system and installed it on the AdhikariLab computer named 'django'. It worked immediately.

I modified the main .m file to include the 40m's NDS server. When you run it you have to include the path to the NDS
client written by Ben Johnson.

The attached is a screenshot of it working on a Mac; it looks as cool on Linux.

Its installed in /cvs/cds/caltech/apps/ligoDV/. In matlab you navigate to that directory and then
type addpath('/cvs/cds/caltech/apps/linux/UNIX_NDS_Client_beta2/') to add the NDS client.
On the Solaris machines, type type addpath('/cvs/cds/caltech/apps/solaris9/UNIX_NDS_Client_beta2/') instead.

Then type ligoDV to start it up. Then click away and have fun.

In the example I've selected
C1:PEM-BS_ACC_EAST_Z
and plotted its specgram.

Big grin


Download and installation instructions, as well as a few examples for use
can be found here (typical lsc username and password):

https://www.gravity.phy.syr.edu/dokuwiki/doku.php?id=ligodv:home
https://www.gravity.phy.syr.edu/dokuwiki/doku.php?id=ligodv:downloading_the_ligodv_software
  143   Thu Nov 29 19:35:14 2007 ranaHowToComputer Scripts / ProgramsGPIB Scripts

Quote:
I've spent a lot of time trying to configurate the GPIB-USB interface for the HP4195. After installing 1) the Agilent libraries, 2) the drivers, 3) the matlab Instrument Toolbox, 4) Jamie script, 5) Alice's script the computer can see the HP but still they can't 'talk' to each other.
I give up. I asked Alice Wang how she managed to get data. I'm not sure she used the GPIB interace. Rob said she might have used the old fashion floppy disks that we can't read anymore here.
I would really appreciate any suggestion by anyone who happened to have the same problems.


Alice and Jamie used the USB-GPIB interface. You should just try using the black laptop which already has this capability or ask Jamie Rollins
who actually knows something.
  142   Thu Nov 29 18:10:13 2007 AlbertoHowToComputer Scripts / ProgramsGPIB Scripts
I've spent a lot of time trying to configurate the GPIB-USB interface for the HP4195. After installing 1) the Agilent libraries, 2) the drivers, 3) the matlab Instrument Toolbox, 4) Jamie script, 5) Alice's script the computer can see the HP but still they can't 'talk' to each other.
I give up. I asked Alice Wang how she managed to get data. I'm not sure she used the GPIB interace. Rob said she might have used the old fashion floppy disks that we can't read anymore here.
I would really appreciate any suggestion by anyone who happened to have the same problems.
  141   Thu Nov 29 15:17:53 2007 robConfigurationPSLISS

I put some ISS beam on the diode on the PSL table. In the previous layout, this was the monitor diode (and it's labeled monitor) but I plugged it into the sensor jack anyways so we can run with the loop closed for now; we can just switch the cables later. The reason the beam was unclear is because someone popped up a flipper mirror which redirects the beam from the ISS into an OSA.

With the ISS gain slider at 15 dB the UGF is around 40kHz.

Why do we have such short cables for the ISS diodes?
  140   Thu Nov 29 14:29:22 2007 tobinConfigurationComputerslinux1 httpd/conlogger fixed
I think I fixed the conlogger web interface on linux1.

Steps necessary to do this:
0. Run "/etc/init.d/httpd start" to start up httpd right now
1. Run "/usr/sbin/ntsysv" and configure httpd to be started automatically in the future
2. Copy /cvs/cds/caltech/conlogger/bin/conlog_web.pl to /var/www/cgi-bin and chown to controls
8. Hack the conlog_web.pl to (0) use /usr/bin/perl (1) not use Apache::Util, and (2) function with the newer version of CGI.pm
9. Enjoy!

The following steps are optional, and may be inserted between steps 2 and 8:
3. Try to install Apache::Util (via "perl -MCPAN -e shell" followed by "Install Apache::Util")
4. Notice that the installation dies because there is no C compiler installed
5. Bang head in disgust and abomination over a Linux distribution shipping without a C compiler installed by default
6. "yum install gcc"
7. Annoyed by further dependencies, go to step 8
  139   Thu Nov 29 11:10:54 2007 robOmnistructureVACRGAlogger sleeping
Without the RGA controller responding, the RGAlogger script just hangs. Rather than fix it, I just put it to sleep by commenting out the line in op440m crontab file. Once we get it running again, we'll move the cronjob to op340m.
  138   Thu Nov 29 10:36:47 2007 albertoConfigurationComputer Scripts / ProgramsAgilent 82357B GPIB to USB Interface Installation Procedeure
To run the Agilent Automation-Ready CD provided with the interface is only the first step of the installation. Apparently there should be also a second CD with the drivers for Windows XP but I couldn't find it. So, after Installaing the IO Libraries Suite from the CD, I had to install the drivers with an executable downloaded from the Agilent's website at:

http://www.home.agilent.com/agilent/editorial.jspx?cc=US&lc=eng&ckey=1188958&nid=-35199.0.00&id=1188958

and only then I could plug in the interface.
Anyway, I burned a cd with the file and put it together with the other one.
  137   Wed Nov 28 21:51:52 2007 tobinConfigurationPSLISS
I replaced the front-end differential receivers for the ISS's "inner-loop" sensor and monitor diode inputs with lower-noise THS4131's (formerly THS4151's). I verified operation by taking the transfer function from the "PD+" and "PD-" inputs (separately) to the testpoint following the differential receiver; the surgery appears successful.

I measured the dark spectra at the ISS's DC PD BNC ports and found a noise floor of ~ 16 nV/rtHz, compared with a floor of ~ 22 nV/rtHz last week. This seems to add up, assuming the DC PD port has 0dB gain: the 4131 has a rated noise of 1.3 nV/rtHz and the 4151 a noise floor of 7.6 nV/rtHz, a difference of 6 nV/rtHz. The other change made in that time was to add a larger power supply bypass capacitor in the PD.

There are two of the old 4151 chips still on the ISS board on the two "outer-loop" channels that we don't use. If I dig up any more 5131's I will replace these too for completeness.

There is currently no light on the ISS diodes; I'm not sure where it's intended to come from.
  136   Wed Nov 28 19:44:18 2007 tobinUpdatePSLHEPA
I found the HEPA turned off completely. I turned it on.
  135   Wed Nov 28 19:02:41 2007 AndreyBureaucracyWIKI-40M UpdateNew WIKI-40M page describing Matlab Suspension Modeling

I created the WIKI-40m page with some details about my today's talk on the 40-m lab meeting.

The address is:

http://lhocds.ligo-wa.caltech.edu:8000/40m/Modeling_of_suspensions

(or you can go to the main page, http://lhocds.ligo-wa.caltech.edu:8000/40m/ , and click on the link "Modeling of suspensions").

The WIKI-40m page describes my transfer functions and contains the pdf-file of my presentation.
  134   Wed Nov 28 17:41:34 2007 robUpdatePSLFSS again
I investigated the FSS a bit more today. I looked at the signals coming out of the FSS frequency reference, and saw that both the LO and PC drive were distorted, non-symmetric waveforms. In addition, the LO path had a 3dB attenuator, meaning the mixer was starved. I placed mini-circuits SLP-30 filters in both paths, and now both are nice sine waves. I also took out the 3dB att. With this work, and the CG slider maxed out at 30, the FSS open loop gain (for real this time) goes up to ~250kHz. Still needs more investigation.
  133   Wed Nov 28 17:15:26 2007 ranaConfigurationSUSETMY damping / watchdogs
Steve has noted that ETMY was often tripping its watchdog. I saw this again today.

So I checked the damping settings. Someone had set the SIDE gain to +1. The gain which gives
it a Q of ~10 is +10. I set the SIDE gain to +20. I checked and the ETMX gain is -16 so now
they're at least similar. I have updated the snapshot to reflect the new value.

Hopefully now it will be more well behaved.
  132   Wed Nov 28 16:46:28 2007 ranaConfigurationComputersscientific linux 5.0
I tried installing Scientific Linux on Tiramisu. The installation process was so bad (really)
that I quit after 15 minutes. Its back to booting Ubuntu as if nothing had ever happened. Let
us never speak of Scientific Linux again.
  131   Wed Nov 28 16:18:15 2007 AlbertoMetaphysicsEnvironmentso clean you can eat on it
I tidied up the desks in the lab, brought the Spectrum Analyzers back to the Salumeria (you don't want to know about that), sorted a lot of stuff and boxed up what I didn't know (you can find it in a couple of carton boxes on the table).
The blackmail with the pie might not work next time.
Please, preserve the common sort.


Alberto
Attachment 1: DSC_0180.JPG
DSC_0180.JPG
Attachment 2: DSC_0181.JPG
DSC_0181.JPG
  130   Wed Nov 28 12:43:53 2007 AndreyBureaucracy Here was the PDF-file of my presentation

I was making a report with powerpoint presentation during that Wednesday's 40-m meeting.

Here was the pdf-file, but LATER IN THE EVENING I CREATED A WIKI-40M-page describing the algorithm, and now the pdf-file is ON THAT WIKI-40M PAGE.


NOTE ADDED AFTER THE PRESENTATION: I double checked, I am indeed taking the root-mean-square of a difference, as we discussed during my talk.

My slide #17 "Calculation of differential length" was wrong, but now I corrected it.
  129   Wed Nov 28 08:47:29 2007 steveOmnistructureVACrga is out of order
The rga is not working since Nov 10
The controller is broken.
pd65-m-d23
Attachment 1: pd65d23.jpg
pd65d23.jpg
  128   Wed Nov 28 04:21:46 2007 ranaUpdatePSLFSS

Quote:
Rana, Tobin

We looked at the RF PD signal to the FSS (siphoning off a signal via a minicircuits directional coupler) and also took an open loop transfer function of the FSS. In the transfer function we saw the step at 100 kHz (mentioned by Rob) as well as some peculiar behavior at high frequency. The high frequency behavior (with a coupling of ~ -20 dB) turns out to be bogus, as it is still present even with the beam blocked. Rearranging the cabling had no effect; the cause is apparently inside the FSS. The step at 100 kHz turns out to be a saturation effect, as it moved as we lowered the signal amplitude, disappearing as we approached -60 dBm. (Above the step, the measurement data is valid; below, bogus.)

Transfer functions will be attached to this entry.

Some things to check tomorrow: the RF signal to the PC, RF AM generation by the PC, LO drive level into the FSS, RF reflection from the PC, efficiency of FSS optical path, quality of RF cabling.


I would also add to Tobin's entry that we believe what Rob was seeing was saturation.

With the bi-directional coupler in there, the RF signal into the FSS board clearly went UP if moved the offset slider away from zero.
With a scope looking at the IN2 testpoint, we can see that there's less than 2 mV offset at zero slider offset.

One tangential thing we noticed with the coupler is that, in lock, the amount of reflected RF is around the same as that going in to the mixer.
I have always wanted to look at this but have only had uni-directional couplers in the past. I think that the double balanced mixer is inherently
not a 50 Ohm device during the times where the diodes are being switched. IF that's the case we might do better in the future by having an RF
buffer on board just before the mixer to isolate the PD head from these reflections.
  127   Tue Nov 27 20:47:00 2007 tobinUpdatePSLFSS
Rana, Tobin

We looked at the RF PD signal to the FSS (siphoning off a signal via a minicircuits directional coupler) and also took an open loop transfer function of the FSS. In the transfer function we saw the step at 100 kHz (mentioned by Rob) as well as some peculiar behavior at high frequency. The high frequency behavior (with a coupling of ~ -20 dB) turns out to be bogus, as it is still present even with the beam blocked. Rearranging the cabling had no effect; the cause is apparently inside the FSS. The step at 100 kHz turns out to be a saturation effect, as it moved as we lowered the signal amplitude, disappearing as we approached -60 dBm. (Above the step, the measurement data is valid; below, bogus.)

Transfer functions will be attached to this entry.

Some things to check tomorrow: the RF signal to the PC, RF AM generation by the PC, LO drive level into the FSS, RF reflection from the PC, efficiency of FSS optical path, quality of RF cabling.
Attachment 1: fss-tf0001.pdf
fss-tf0001.pdf fss-tf0001.pdf
  126   Tue Nov 27 16:18:58 2007 robConfigurationIOOMC loop
Reduced the common gain to 22dB in the mcup script, so that the WFS would not blow the lock. The above measure of the OLG was done without the mcWFS running, so may be a low estimate as compared to when the alignment is perfect.
  125   Tue Nov 27 15:47:17 2007 robConfigurationIOOMC loop
After the FSS running pretty quick, I checked the MC loop. I used TPA 1&2.

MC loop
UGF: 70kHz
Input Gain: 29dB
Boost Level: 2
phase: 40 deg
Attachment 1: MCsmall.jpg
MCsmall.jpg
  124   Tue Nov 27 15:45:08 2007 robConfigurationPSLFSS loop

It's unclear (to me, at least) what was the end result of the FSS path tweaking before Thanksgiving. Today I measured the open loop gain, and it was still around 100kHz, even with the gain sliders maxed out, but it looked really crappy with a sharp cutoff around the UGF. Then, on a lark, I pushed around the "Input Offset Adjust" slider, which sums an offset into the signal coming out of the mixer. By moving this slider to 7V, I got the UGF to 500kHz with 45 deg of phase. That would be fine, and we could go offset hunting, but the same thing happens if one puts in a large negative value! I don't really understand what's going on, but it seems like weirdness in the electronics. Unfortunately the web interface to the conlog is not running (presumably because the `new' linux1 doesn't have its apache server running) and my command line conlog efforts have been stymied. So, I don't know what the historical settings of this offset are, but zero is definitely not a good setting right now. Here's a snapshot:

FSS
UGF: 500kHz
CG : 24dB
FG : 19dB
input offset: 7V
Phase Adjust: 1.09V
Phase Button: 0
RF Amp Adjust: 7.38V

margins:
phase: 45 deg
gain: 8dB
Attachment 1: FSSsmall.jpg
FSSsmall.jpg
  122   Mon Nov 26 10:17:31 2007 steveOmnistructureSUSetmy sus damping restored
20 days plot is showing etmy loosing damping 4 times.
I zoomed in with each event. Three of them could of been triggered
by garbage loading just outside. However attachment 2 plot demonstrating that small earthquake or seismic event
did not tripped etmy damping.
The fourth event was preceded by a 4-5 hrs of continous rise of the rms motion at C1:SUS-ETMY_LLPD_VAR
Attachment 1: etmyrms20d.jpg
etmyrms20d.jpg
Attachment 2: etmyrmseq.jpg
etmyrmseq.jpg
  121   Wed Nov 21 14:31:41 2007 robUpdatePSLFSS twiddle

I `tweaked' the FSS path today. Here's what I did:

1) Shut down the FSS autolocker

2) Turn off FSS servo

3) Assume the beam coming back from the AOM is double-first-order, and don't make any changes large enough to lose it.

4) Tweak the alignment of these components to maximize the incident power on the RC reflected diode:

a) PBS before AOM
b) AOM
c) curved mirror after the AOM

5) Translate the AOM such that the beam moves away from the PZT, then when it levels off (no more power gains with movement),
move it back just a little bit so there's a teensy drop in power. This should but the beam as close to the edge as possible,
but whether or not it's the best place is still to be determined.

6) Lock the FSS, and align the mirrors into the frequency reference cavity.

After all this, the RC transmitted power went from .57 to .73 -- probably not a big enough change to account for the missing loop
gain, but we'll know more once the loop gets measured (after Alberto stops hogging the Agilent network analyzer).

Other possible routes include a systematic check of the upstream path (e.g., the Pockels cell) and just increasing the pickoff fraction for the FSS.
  120   Tue Nov 20 18:35:20 2007 JohnHowToComputersMatLab in Emacs
If you can't get MatLab to run in emacs try adding the following to the .emacs file

(setq matlab-shell-command-switches '("-nojvm"))

This stops the gui opening.

To start MatLab type M-x matlab-shell.
To enter MatLab mode M-x matlab-mode.

I've done this on LINUX3.

To run MatLab in emacs under windows one can use MatLabShell http://www.cs.umb.edu/~ram/matlabShell/index.html
  119   Tue Nov 20 18:02:54 2007 JohnSummaryComputersPSL_Main screen
I've updated the PSL_MAIN screen. The old version may be found in cvs/cds/caltech/medm/old/medm/psl.
Attachment 1: PSL_Screen.tif
PSL_Screen.tif
  118   Tue Nov 20 13:06:57 2007 tobinConfigurationComputerslinux1 has new disk
Alex put the new hard disk into linux1 along with a fresh install of linux (CentOS). The old disk was too damaged to copy.

Alex speculates that the old disk failed due to overheating and that linux1 could use an extra fan to prevent this in the future.
  117   Tue Nov 20 11:10:07 2007 tobinUpdateComputersepics access from matlab
I installed "labca", which allows direct access to EPICS channels from within Matlab. It comes with both Linux and Solaris binaries (and source) but I've only tried it on linux.

To set it up, run these shell commands:
pushd /cvs/cds/caltech/users/tf/build/labca_2_1/bin/linux-x86
setenv PATH ${PATH}:`pwd`
cd /cvs/cds/caltech/users/tf/build/labca_2_1/lib/linux-x86
setenv LD_LIBRARY_PATH ${LD_LIBRARY_PATH}:`pwd`
popd
Then start matlab, and within matlab type:
addpath /cvs/cds/caltech/users/tf/build/labca_2_1/bin/linux-x86/labca
help labca
foo = lcaGet('C1:PSL-FSS_RCTRANSPD')
It seems like reasonably well-written software, and is being actively maintained right now. If we like it, I can build a more recent version, install it in a more permanent location, etc.
  116   Tue Nov 20 10:11:33 2007 JohnSummaryPSLPMC pole measurements
We measured the PMC pole in the following way.

1. Reduced laser power by rotating lambda/2 plate at laser output. Thermal effects in the PZT distort resonance peaks. Reducing power too much leads to problems with digitisation error.

2. Sweep NPRO PZT (C1: PSL-FSS_INOFFSET) using trianglewave. Record ramp, PMC transmission and reference cavity transmission ('C1: PSL-FSS_FAST','C1: PSL-ISS_INMONPD_F','C1: PSL-FSS_RCTRANSPD_F).

3. Since the PZT cannot sweep a full FSR in the PMC we looked at the sideband resonances within the reference cavity to calibrate the actuator.
Result: 7.35 +/- 0.22 MHz/V

4. Use #3 to calibrate the x axis of the PMC transmission.

5. Fit PMC resoances to an Airy function to get finesse. Take an average, weighted according to the resnorm. Calculate cavity pole frequency.
Result: 380kHz +/- 59kHz. This corresponds to a finesse of ~936. According to this plot the nominal pole is at 488kHz and the finesse is 732.

This is by no means a definitive measurement due to the misshapen resonance peaks recorded.
Attachment 1: FittedPMCPeak.jpg
FittedPMCPeak.jpg
  115   Mon Nov 19 14:32:10 2007 steveBureaucracySAFETYgrad student safety training
John Miller and Alberto Stochino has received the 40m safety bible.
They still have to read the laser operation manual and sign off on it.
  114   Mon Nov 19 14:19:25 2007 steveOmnistructurePEMjackhammer
The construction personal successfully jackhemmered a fence around the "Drever's parking slot"
There is no parking space available close by
Attachment 1: jackhammer.jpg
jackhammer.jpg
Attachment 2: jackhammer2.jpg
jackhammer2.jpg
  113   Fri Nov 16 18:46:49 2007 steveBureaucracyPSL MOPA was turned off & on
The "Mohana" boys scouts and their parents visited the 40m lab today.
The laser was turned off for their safety.
It is back on !
  112   Fri Nov 16 14:31:43 2007 tobinUpdateComputersop140 disks
Phil Ehrens stopped by and took op140's disks.
Attachment 1: DSC_0173.JPG
DSC_0173.JPG
  111   Fri Nov 16 14:11:26 2007 tobinUpdateComputersop140
Alan called to say that Phil Ehrens will be coming by to take op140 off our hands.
  110   Fri Nov 16 11:27:18 2007 tobinUpdateComputersscript fix
I added a tidbit of code to "LIGOio.pm" that fixes a problem with ezcastep on Linux. Scripts such as "trianglewave" will now work on Linux.
# On Linux, "ezcastep" will interpret negative steps as command line arguments,
# because the GNU library interprets anything starting with a dash as a flag.
# There are two ways around this.  One is to set the environment variable
# POSIXLY_CORRECT and the other is to inject "--" as a command line argument
# before any dashed arguments you don't want interpreted as a flag.  The former
# is easiest to use here:

if (`uname` =~ m/Linux/) {
    # Add an environment variable for child processes
    $ENV{'POSIXLY_CORRECT'} = 1;
}
  109   Thu Nov 15 18:37:06 2007 tobinUpdateComputerspossible replacement for linux1's disk
It looks like the existing disk in linux1 is a Seagate ST380013A (this can be found either via the smartctl utility or by looking at the file /proc/ide/hda/model). It appears that you can still buy this disk from amazon, though I think just about any ATA disk would work. I'll ask Steve to buy one for us.
  108   Thu Nov 15 18:36:48 2007 JohnSummary PSL table work
I've rotated the lambda/2 plate to 340deg (from 6 deg) and blocked one arm of the Mach-Zender. Undo both if you need to.
  107   Thu Nov 15 18:23:55 2007 JohnHowToComputersSwap CAPS and CTRL on a Windows 2000/XP machine
I've swapped ctrl and caps on the four control room Windows machines. Right ctrl is unchanged.



Start menu->Run "regedit"

HKEY_LOCAL_MACHINE\SYSTEM\CurrentControlSet\Control\Keyboard Layout

Click on the KeyboardLayout entry.

Edit->New Binary Value name it Scancode Map.

Then select the new Scancode Map entry.

Edit menu->Modify Binary Data.

In the dialog box enter the following data:

0000: 00 00 00 00 00 00 00 00
0008: 03 00 00 00 3A 00 1D 00
0010: 1D 00 3A 00 00 00 00 00

Exit the Registry Editor. You need to log off and then on in XP (and restart in Windows 2000) for the changes to be made.
  106   Thu Nov 15 18:06:06 2007 tobinUpdateComputersalex: linux1 root file system hard disk's dying
I just noticed that Alex made an entry in the old ilog yesterday, saying: "Looks like linux1 root filesystem hard drive is about to die. The system log is full of drive seek errors. We should get a replacement IDE drive as soon as possible or else the unthinkable could happen. 40 Gb IDE hard drive will be sufficient."
  105   Thu Nov 15 17:09:37 2007 pkpUpdateOMCVertical Transfer functions with no cables attached.
[Norna Pinkesh]

The cables connecting all the electronics ( DCPDs, QPDs etc) have been removed to test for the vertical transfer function. Now the cables are sitting on the OMC bench and it was realigned.
Attachment 1: VerticaltransferfuncnocablesattachedNov152007.pdf
VerticaltransferfuncnocablesattachedNov152007.pdf VerticaltransferfuncnocablesattachedNov152007.pdf VerticaltransferfuncnocablesattachedNov152007.pdf VerticaltransferfuncnocablesattachedNov152007.pdf
  104   Thu Nov 15 04:18:11 2007 JohnSummaryPSLPMC cavity pole measurements
In connection with our work on the ISS I attempted to measure the PMC cavity pole.

I swept the PMC PZT and looked at the transmission through the cavity on the ISS Monitor diode (which is now back on the table, feel free to remove it again tomorrow).

To avoid thermal effects I reduced the laser power using the half wave plate at the laser ouput (rotated from 6 deg to 340deg).

I swept the PZT using the triangle wave command "trianglewave C1: PSL-PMC_RAMP -3.5 3.3 20 200". I noticed that the functional form of the resonances deteriorated over the duration of the excitation. Each sweep was able to capture just over one FSR. The resonances were a little close to the 'points' of the triangle wave for my liking although I don't think PZT hysteresis was a big factor.

Looking at the data the peaks are not of uniform width across a sweep or between consecutive sweeps. Hence any results from this mesurement are not particularly useful. I can't be sure if this was due to misalignments, thermal effects, higher order mode content or some other affect.

Rob suggests sweeping the laser frequency using the NPRO PZT instead.
Attachment 1: Peaks.jpg
Peaks.jpg
  103   Wed Nov 14 17:50:00 2007 tobinUpdatePSLISS
Here's the current wiring between the ISS and its PDs:

pin cable PD ISS
1 blue +5 +5
2 red +15 +15
3 white -15 -15
4 brown OUT IN PD +
5,6,7,8 no connection no connection GND
9 black GND IN PD -


The schematics for the ISS and the PDs are linked from our wiki.

We'll connect the ISS GND to the PD GND.
  102   Wed Nov 14 16:54:54 2007 pkpUpdateOMCMuch better looking vertical transfer functions
[Norna Pinkesh]

So after Chub did his wonderful mini-surgery and removed the peek from the cables and after Norna and I aligned the whole apparatus, the following are the peaks that we see.
It almost exactly matches Norna's simulations and some of the extra peaks are possibly due to us exciting the Roll/longitudnal/yaw and pitch motions. The roll resonance is esp prominent.

We also took another plot with one of the wires removed and will wait on Chub before we remove another wire.
Attachment 1: VerticalTransPreampwireremovedNov142007.pdf
VerticalTransPreampwireremovedNov142007.pdf VerticalTransPreampwireremovedNov142007.pdf VerticalTransPreampwireremovedNov142007.pdf VerticalTransPreampwireremovedNov142007.pdf
Attachment 2: VerticalTranswiresclampedNov142007.pdf
VerticalTranswiresclampedNov142007.pdf VerticalTranswiresclampedNov142007.pdf VerticalTranswiresclampedNov142007.pdf VerticalTranswiresclampedNov142007.pdf
  101   Wed Nov 14 12:47:19 2007 tobinUpdatePSLISS
John, Tobin

With John's notch filter installed and the increased light on the ISS sensing diode, we were able to get a UGF of about 60 kHz with the gain slider set to about 20 dB. This morning we met with Stefan to learn his ISS-fu.

His recommendations for the ISS include:
  • Replace the cables from the board to the front panel connectors if this hasn't already been done.
  • Replace the input opamps with 4131's. Be sure to test both positive and negative input signals.
  • Check that all the compensation capacitors are in place and are 68 pF
  • Make sure all the feedback loops have high frequency rolloff
  • The ISS board reads the PDs differentially; make sure the PD sends differentially.
  • Add a big (ie 10uF tantalum) capacitor to the PD to suppress power supply noise
  • Add bigger power supply bypass caps to the ISS
I just took sensing noise spectra (from the PD DC bnc ports) and then took the photodiodes off the table to check that they have the negative end of the differential line connected to ground. (I placed black metal beam blocks on the table in place of the ISS PD's. Also, from the ISS schematic, it looks like it sends a differential output to the PD DC bnc ports, but we have been plugging them directly into the SR785 (grounding the shield). We should make a little BNC-doodle that separates the signal+shield to go into the A and B inputs on the spectrum analyzer.) Opening up one of the photodiodes, it appears that the negative line of the differential output is not connected. Will continue later this afternoon.
  100   Wed Nov 14 12:33:35 2007 tobinAoGEnvironmentconstruction
The construction crews are running a jack-hammer right outside of the control room.
Attachment 1: DSC_0172.JPG
DSC_0172.JPG
  99   Wed Nov 14 07:48:38 2007 nornaOmnistructureOMCOMC Cable dressing
[Snipped from an email]

1) Last Friday Pinkesh and I set the OMC up with only the top three OSEMs and took a vertical transfer function. We had removed the other OSEMs due to difficulty of aligning all OSEMs with the weight of the bench etc bringing the top mass lower than the tablecloth can accommodate. See attached TF.Clearly there are extra peaks (we only expect two with a zero in between) and my belief is that at least some of them are coupling of other degrees of freedom caused by the electrical wiring. Pinkesh and I also noticed the difficulty of maintaining alignment if cables got touched and moved around. So.....

2) Yesterday Dennis and I took a look at how much moving a cable bundle around (with the peak shielding) changed the DC alignment. In a not too precise experiment ( using HeNe laser reflecting off the bench onto a surface ~ 1 metre away) we saw that we could reposition the beam one or two mm in yaw and pitch. This corresponds to ~ one or two mrad which is ~ the range of the OSEM DC alignment. We discussed possibility of removing the cabling from the middle mass, removing the peak and taking it from the bench directly to the structure above. I asked Chub if he could make an equivalent bundle of wires as those from the two preamps to see what happens if we repeat the "moving bundle" experiment. So...

3) Today Chub removed the cabling going to the preamps and we replaced it with a mock up of wire bundle going directly from the preamps to the structure above. See attached picture. The wires are only attached to the preamp boxes weighted down with masses but the bundle is clamped at the top. We repeated the "wiggle the bundle" test and couldnt see any apparent movement ( so maybe it is at most sub-mm). The cable bundle feels softer.

The next thing Chub did was to remove the second bundle ( from photodiodes, heater, pzt) from its attachment to the middle mass and strip off the peek. It is now also going to the top of the structure directly. The whole suspension now appears freer. We discussed with Dennis the "dressing " of the wires. There are some minor difficulties about how to take wires from the bright side to the dark side, but in general it looks like that the wires forming the second "bundle" could be brought to the "terminal block" mounted on the dark side and from there looped up to the top of the structure. We would have to try all this of course to see the wiring doesnt get in the way of other things (e.g. the L and R OSEMs). However this might be the way forward. So...

4) Tomorrow Pinkesh and I will check the alignment and then repeat the vertical transfer function measurement with the two bundles as they are going from bench to top of structure. We might even do a horizontal one if the middle mass is now within range of the tablecloth.
We can then remove preamp cables completely and lay the second bundle of cables on the optical bench and repeat the TFs.

The next thing will be to weigh the bench plus cables. This will allow us to
a) work out what counterbalance weights are needed - and then get them manufactured
b) firm up on how to handle the extra mass in terms of getting the masses at the correct height.

And in parallel Chub will work on the revised layout of cabling.

Looking a little further ahead we can also get some stiffness measurements made on the revised bundle design ( using Bob's method which Alejandro also used) and fold into Dennis's model to get some sanity check the isolation.

I think that's it for now. Comments etc are of course welcome.

Norna
Attachment 1: OMC-11-13-07_011.jpg
OMC-11-13-07_011.jpg
Attachment 2: VerticalTrans.pdf
VerticalTrans.pdf VerticalTrans.pdf VerticalTrans.pdf VerticalTrans.pdf
  98   Tue Nov 13 14:33:40 2007 JohnUpdatePSLISS filter
The transfer function from 'In Loop Error Point Monitor' to TP3 the filter out test point on the ISS board.

-33dB at 3.715MHz.
Attachment 1: PB130035.JPG
PB130035.JPG
Attachment 2: DSC_0165.JPG
DSC_0165.JPG
  97   Mon Nov 12 23:44:19 2007 JohnUpdatePSLISS

Quote:

After John soldered a 3.7 MHz notch filter onto the ISS board, I took a quick TF and RIN measurement. The out-of-loop RIN is attached, including a dark noise trace, and with the gain slider at 10dB. The UGF is 35kHz with a phase margin of 30deg. John is currently doing a more thorough inspection, and will detail his findings in a subentry.


No progress on the ISS tonight. I tried to implement a new filter (attached)to try and gain some phase before the notch. If anything this made things worse. More work is needed.

The ISS loop is off and the power is off at the chassis.
Attachment 1: ISSfilter.jpg
ISSfilter.jpg
  96   Mon Nov 12 15:18:34 2007 robUpdatePSLISS

After John soldered a 3.7 MHz notch filter onto the ISS board, I took a quick TF and RIN measurement. The out-of-loop RIN is attached, including a dark noise trace, and with the gain slider at 10dB. The UGF is 35kHz with a phase margin of 30deg. John is currently doing a more thorough inspection, and will detail his findings in a subentry.
Attachment 1: ISS.png
ISS.png
ELOG V3.1.3-